# TreeDGS: Aerial Gaussian Splatting for Distant DBH Measurement

### v1

@@ Abstract
Aerial remote sensing enables efficient large-area surveying, but accurate direct object-level measurement remains difficult in complex natural scenes. Recent advancements in 3D vision, particularly learned radiance-field representations such as NeRF and 3D Gaussian Splatting, have begun to raise the ceiling on reconstruction fidelity and densifiable geometry from posed imagery. Nevertheless, direct aerial measurement of important natural attributes such as tree diameter at breast height (DBH) remains challenging. Trunks in aerial forest scans are distant and sparsely observed in image views: at typical operating altitudes, stems may span only a few pixels. With these constraints, conventional reconstruction methods leave breast-height trunk geometry weakly constrained. We present TreeDGS, an aerial image reconstruction method that leverages 3D Gaussian Splatting as a continuous, densifiable scene representation for trunk measurement. After SfM–MVS initialization and Gaussian optimization, we extract a dense point set from the Gaussian field using RaDe-GS’s depth-aware cumulative-opacity integration and associate each sample with a multi-view opacity reliability score. We then estimate DBH from trunk-isolated points using opacity-weighted solid-circle fitting. Evaluated on 10 plots with field-measured DBH, TreeDGS reaches 4.79,cm RMSE (about 2.6 pixels at this GSD) and outperforms a state-of-the-art LiDAR baseline (7.91,cm RMSE), demonstrating that densified splat-based geometry can enable accurate, low-cost aerial DBH measurement (Fig. LABEL:fig1).
@@ 1 Introduction
Aerial remote sensing has quickly become a cornerstone of modern environmental monitoring for its ability to survey large areas rapidly and cost efficiently. [1, 7, 36]. Yet, despite major gains in sensor resolution, platform stability, and reconstruction pipelines, extracting direct, object-level measurements from aerial imagery remains difficult in structurally complex natural scenes [8, 15]. Forested environments are particularly challenging: heterogeneous geometry, self-similar textures, frequent occlusion, and strong appearance variability across lighting conditions all weaken the visual cues that aerial reconstruction methods rely on [15].
A prominent example is tree diameter at breast height (DBH)—a foundational forestry variable that aerial remote sensing struggles to recover reliably. Its importance for forestry cannot be understated: DBH is among the most information-dense descriptors of an individual tree’s size and growth stage. It is a primary input to standard allometric models and reporting workflows that estimate wood volume, biomass, and carbon [2, 17, 5, 18]. These allometric models, in turn, underpin decision making used across forestry operations and policy, ranging from timber inventory and silvicultural planning to carbon offset quantification, wildfire risk mitigation, and long-term ecosystem management [41, 3].
DBH estimation from above-canopy aerial sensing is fundamentally constrained by an observability bottleneck. In typical UAV image surveys flown at operational altitudes, stems are not “large objects.” In our data, captured at $\sim$70 m above ground ($\mathrm{GSD}\approx 1.84$ cm/px), stems in the evaluation plots span only $\sim$15 pixels across per view on average (Table 1), and can be just a few pixels wide for smaller trees (Fig. 2). This pixel scarcity is compounded by view scarcity: the breast-height band ($h_{\mathrm{BH}}\approx 1.37$ m in our field protocol) is frequently occluded by crowns, branches, and understory vegetation, so each stem may be seen only in a handful of oblique glimpses through noisy canopy gaps. The core geometric challenge of measurement is therefore recovering enough faithful breast-height trunk surface to support stable circle or cylinder fitting under a regime of few pixels, few views, and partial visibility.
@@
Recent learned radiance-field representations, such as Neural Radiance Fields (NeRF) [33] and 3D Gaussian Splatting (3DGS) [22], can improve reconstruction fidelity from posed imagery and offer a promising path forward for thin structures that are only weakly observed in individual frames. Early forestry applications have begun exploring radiance fields for forest monitoring and tree-scale reconstruction [14, 23, 42, 27], but the problem of direct DBH measurement from stand-off, above-canopy UAV surveys remains largely unaddressed. Moreover, translating a “fluffy” radiance-field reconstruction into a reliable diameter requires a measurement-oriented extraction step: sufficiently dense sampling at breast height, and a way to prioritize samples that are consistently supported across views. Most forestry applications of radiance fields to date have focused on close-range terrestrial [27] imagery which have a different set of constraints, leaving aerial DBH measurement largely unaddressed by learned radiance field methods.
We introduce TreeDGS, an aerial reconstruction pipeline that uses 3D Gaussian Splatting as a measurement-oriented scene representation for estimating DBH from commodity UAV RGB imagery under pixel- and view-limited trunk observations. Starting from standard SfM camera poses and an MVS initialization, we optimize a RaDe-GS model [50] to obtain a continuous Gaussian field whose covariances and opacities are refined by multi-view photometric consistency. We then extract a dense point set using depth-aware cumulative-opacity integration, retaining samples whose accumulated opacity indicates a reliably occupied surface under a fixed threshold. To mitigate spurious samples from foliage and partial occlusions, we compute a per-point multi-view opacity support score as a reliability signal and retain trunk-isolated points via 3D semantic segmentation. Finally, we estimate DBH with opacity-weighted, slice-wise circle fitting, yielding stable diameter estimates even when the trunk is only weakly observed in individual images.
@@
  • We present TreeDGS, a method that directly reconstructs and measures DBH from above-canopy UAV RGB imagery by leveraging a 3D Gaussian Splatting scene representation in a pixel-limited regime.
@@
The remainder of the paper describes our data collection and evaluation protocol (Sec. 2), details the proposed pipeline and opacity-based sampling strategy (Sec. 3), and presents experimental results and ablations (Sec. 4).
@@ 2 Materials
Data were collected in a managed Pinus taeda (loblolly pine) stand in southeastern Arkansas, USA. We established ten 0.2-acre circular plots (plots 1–10; radius 16.05 m; Fig. 3).
At each plot center, we marked a circular plot boundary with a 16.05 m radius and measured all stems within the boundary. Each included tree was assigned a unique ID. For each tree, we recorded: (i) bearing and distance from the plot center; (ii) DBH measured with a diameter tape at 1.37 m above ground level; and (iii) height measured using a laser rangefinder.
@@
For the full set of plots, we collected overlapping RGB imagery using a DJI Matrice 4E in two campaigns. Imagery was captured with high overlap and mixed viewing geometry (including oblique views at 70 and 90 m altitude) to support SfM/MVS reconstruction and increase the chance of observing trunk surfaces through gaps in the canopy and understory. To improve absolute positioning consistency across flights, we corrected the onboard UAV GNSS trajectory via post-processed kinematic (PPK) processing using reference data from a nearby CORS station (ARMO), and used the corrected solution to update image geotags before reconstruction. An example UAV RGB flight trajectory is shown in Fig. 4(a).
Figure 5: TreeDGS pipeline. RGB images are reconstructed with SfM/MVS and represented as 3D Gaussian splats $G$. We surface-sample dense points using opacity $\alpha$ and weights $w$, segment trunks $T$, and estimate DBH by opacity-weighted circle fitting on multiple trunk slices.
@@ 3.1 Problem Statement
where $\mathbf{P}_{i}$ are the calibrated camera matrices, $\mathcal{X}_{\text{dense}}$ is a densified point set from OpenMVS, $\mathcal{G}$ is the optimized Gaussian field, and $\mathcal{T}_{t}$ denotes trunk points for tree instance $t$. Sections 3.23.6 detail SfM/OpenMVS reconstruction, Gaussian optimization, trunk extraction through 3D segmentation [48] with opacity cues, and the final DBH fitting procedure.
@@ 3.2 Structure-from-Motion and Multi-View Stereo
SfM takes the matched correspondences across overlapping image pairs and estimates: (i) camera poses $\{\mathbf{P}_{i}\}$ and (ii) a sparse 3D point set. We use standard robust estimation and bundle adjustment from GLOMAP [37] to refine the reconstruction. We keep notation minimal and denote the resulting calibrated cameras as:
@@
These cameras are used both for Gaussian training (Section 3.3) and for the multi-view opacity tests during surface sampling (Section 3.4).
@@
Figure 6: Comparison of back-projection points with multi-view depth consistency and the proposed sample points. (a) Depth-fusion baseline: rendered depth points are back-projected into 3D and kept only if they pass multi-view depth-consistency checks. This produces large sheet-like artifacts near the ground/understory and incomplete stem surfaces (red dashed boxes), despite appearing consistent across views [50]. (b) Our opacity-guided surface sampling with depth-aware point-wise compositing [50] yields a cleaner, denser stem geometry suitable for DBH estimation.
@@ 3.3 Reconstruction with Gaussian Splats
Given calibrated cameras $\{\mathbf{P}_{i}\}$ and an MVS-densified initialization $\mathcal{X}_{\text{dense}}$ from Section 3.2, we optimize a 3D Gaussian Splatting (3DGS) scene representation. We adopt RaDe-GS [50], which builds on the real-time 3DGS renderer [22] while explicitly rasterizing depth in a way that is also useful for our subsequent surface-consistent point sampling.
@@
We integrate RLGS [28] as an online controller that adapts selected training hyper-parameters based on observed optimization dynamics. In our setting, this is primarily used to stabilize opacity and geometry quality for the downstream DBH pipeline. After convergence, the optimized Gaussian field is denoted $\mathcal{G}$ and serves as the source representation for dense point sampling.
@@ 3.4 Surface Sampling with Opacity
Our downstream stages (3D trunk segmentation and DBH fitting) require a point cloud that is (i) dense enough to represent thin stems at stand-off distance and (ii) geometrically trustworthy under heavy occlusion from foliage and understory. This is not automatic from a 3D Gaussian Splatting (3DGS) reconstruction: the optimized field $\mathcal{G}$ is a collection of overlapping volumetric primitives whose parameters are trained through front-to-back alpha compositing rather than an explicit surface loss [22, 50]. As a result, two seemingly straightforward exports are unreliable in our UAV forest regime. First, exporting one point per Gaussian (the means) produces a fragmented and extremely sparse cloud (see Section 4 Fig. 10), which is insufficient for robust 3D segmentation and measurement. Second, back-projecting rendered depth maps into 3D and fusing them via multi-view depth-consistency tends to preserve view-consistent sheet-like artifacts from semi-transparent vegetation while eroding already-scarce trunk evidence, especially when stems are only a few pixels wide or intermittently visible (see Fig. 6). We therefore sample densely in 3D, but we accept and score samples using tests that are consistent with the same compositing model that governs 3DGS rendering and optimization.
@@
The per-pixel accumulated alpha mask $A_{v}(\mathbf{u})$ tells us whether a ray intersects reconstructed mass somewhere along the ray, but it does not tell us whether a particular 3D sample $\mathbf{p}$ lies on the visible surface, behind it, or in front of it (all can project to the same pixel). For surface extraction we therefore need a depth-aware test that measures how much opacity accumulates in front of $\mathbf{p}$ along the viewing ray, using the same compositing model that governs training. RaDe-GS provides exactly this query by integrating point-wise alphas with a depth-plane formulation.
@@
The $\min(\cdot)$ depth clamping in Eq. (6) is critical for surface consistency: once $\mathbf{p}$ lies behind the locally rasterized surface depth $d_{k,v}(\cdot)$, the depth residual term saturates rather than increasing with $r_{v}(\mathbf{p})$, preventing spurious “volumetric” accumulation behind the surface. RaDe-GS then composites these point-wise alphas using the same front-to-back transmittance update as in Eq. (5), yielding the cumulative opacity
@@
We then store two per-point reliability signals:
@@ 3.5 Semantic Trunk Extraction
The sampled point cloud from Section 3.4 contains trunks mixed with foliage, understory vegetation, and occasional floating samples caused by Gaussians with nonzero opacity in free space. To isolate stem geometry for DBH measurement, we used ForestFormer3D [48], a 3D semantic segmentation model, to predict a class label for each sampled point and to assign per-tree instance IDs when available. In our pipeline, the per-point opacity reliability $\bar{\alpha}$ is kept as an auxiliary output feature alongside $(x,y,z)$. We retain only points predicted as trunk to form a trunk-only cloud for each tree instance $t$, denoted $\mathcal{T}_{t}=\{(\mathbf{x}_{k},\bar{\alpha}_{k})\}$. This separation step reduces contamination from branches and understory, while the retained opacity values enable reliability-weighted fitting downstream. The trunk-only instances are then passed to the geometric measurement stage described next.
@@ 3.6 Opacity-Weighted DBH Measurement
Even after per-slice robust fitting, some slices can be corrupted by branch attachments, residual foliage, or incomplete sampling. We therefore fit a simple taper model on the lower stem using RANSAC with a negative-slope prior:
@@
Table 2: Per-plot DBH accuracy and robustness (with pooled results). RMSE/MAE/ME are in cm and RRMSE is in %. SR reports successful estimates over total trees. Bold indicates the best within each plot. TreeDGS with opacity-weighted circle fitting attains the best RMSE in 7/10 plots and the best pooled performance.

### v2

@@ Abstract
Aerial remote sensing enables efficient large-area surveying, but accurate direct object-level measurement remains difficult in complex natural scenes. Recent advancements in 3D vision, particularly learned radiance-field representations such as NeRF and 3D Gaussian Splatting, have begun to raise the ceiling on reconstruction fidelity and densifiable geometry from posed imagery. Nevertheless, direct aerial measurement of important natural attributes such as tree diameter at breast height (DBH) remains challenging. Trunks in aerial forest scans are distant and sparsely observed in image views: at typical operating altitudes, stems may span only a few pixels. With these constraints, conventional reconstruction methods leave breast-height trunk geometry weakly constrained. We present TreeDGS, an aerial image reconstruction method that leverages 3D Gaussian Splatting as a continuous, densifiable scene representation for trunk measurement. After SfM–MVS initialization and Gaussian optimization, we extract a dense point set from the Gaussian field using RaDe-GS’s depth-aware cumulative-opacity integration and associate each sample with a multi-view opacity reliability score. Then, we estimate DBH from trunk-isolated points using opacity-weighted solid-circle fitting. Evaluated on 10 plots with field-measured DBH, TreeDGS reaches 4.79,cm RMSE (about 2.6 pixels at this GSD) and outperforms a state-of-the-art LiDAR baseline (7.91,cm RMSE). This shows that TreeDGS can enable accurate, low-cost aerial DBH measurement (Fig. LABEL:fig1).
@@ 1 Introduction
Aerial remote sensing has quickly become a cornerstone of modern environmental monitoring for its ability to survey large areas rapidly and cost efficiently [1, 7, 36]. Yet, despite major gains in sensor resolution, platform stability, and reconstruction pipelines, extracting direct, object-level measurements from aerial imagery remains difficult in structurally complex natural scenes [8, 15]. Forested environments are challenging: heterogeneous geometry, self-similar textures, frequent occlusion, and strong appearance variability across lighting conditions all weaken the visual cues that aerial reconstruction methods rely upon [15].
A prominent example is tree diameter at breast height (DBH)—a foundational forestry variable that aerial remote sensing struggles to recover reliably. Its importance for forestry cannot be understated: DBH measures an individual tree’s size and growth stage, and is a primary input to standard allometric models and reporting workflows that estimate wood volume, biomass, and carbon [2, 17, 5, 18]. In turn, these allometric models underpin decision making used across forestry operations and policy, ranging from timber inventory and silvicultural planning to carbon offset quantification, wildfire risk mitigation, and long-term ecosystem management [41, 3].
DBH estimation from above-canopy aerial sensing is fundamentally constrained by observability. In typical UAV image surveys flown at operational altitudes, stems are not “large objects.” In our data, captured at $\sim$70 m above ground ($\mathrm{GSD}\approx 1.84$ cm/px), stems in the evaluation plots span only $\sim$15 pixels across per view on average (Tab. 1), and can be just a few pixels wide for smaller trees (Fig. 2). This pixel scarcity is compounded by view scarcity: the breast-height band ($h_{\mathrm{BH}}\approx 1.37$ m in our field protocol) is frequently occluded by crowns, branches, and understory vegetation, so each stem may be seen only in a handful of oblique glimpses through noisy canopy gaps. Therefore, the core geometric challenge of measurement is recovering enough faithful breast-height trunk surface to support stable circle or cylinder fitting under a regime of few pixels, few views, and partial visibility.
@@
Recent radiance-field representations, such as Neural Radiance Fields (NeRF) [33] and 3D Gaussian Splatting (3DGS) [22], can improve reconstruction fidelity from posed imagery and offer a promising path forward for thin structures that are only weakly observed in individual frames. Early forestry applications have begun exploring radiance fields for forest monitoring and tree-scale reconstruction [14, 23, 42, 27], but the problem of direct DBH measurement from stand-off, above-canopy UAV surveys remains largely unaddressed. Moreover, translating a “fluffy” radiance-field reconstruction into a reliable diameter requires a measurement-oriented extraction step: a sufficiently dense sampling at breast height and a way to prioritize samples that are consistently supported across views. Most forestry applications of radiance fields to date have focused on close-range terrestrial [27] imagery which have a different set of constraints, leaving aerial DBH measurement largely unaddressed by radiance field methods.
We introduce TreeDGS: an aerial reconstruction pipeline that uses 3D Gaussian Splatting as a measurement-oriented scene representation for estimating DBH from commodity UAV RGB imagery under pixel- and view-limited trunk observations. Starting from standard SfM camera poses and an MVS initialization, we optimize a RaDe-GS model [50] to obtain a continuous Gaussian field whose covariances and opacities are refined by multi-view photometric consistency. Then, we extract a dense point set using depth-aware cumulative-opacity integration, retaining samples whose accumulated opacity indicates a reliably occupied surface under a fixed threshold. To mitigate spurious samples from foliage and partial occlusions, we compute a per-point multi-view opacity support score as a reliability signal and retain trunk-isolated points via 3D semantic segmentation. Finally, we estimate DBH with opacity-weighted, slice-wise circle fitting, yielding diameter estimates even when the trunk is only weakly observed in individual images.
@@
  • We present TreeDGS, a method that directly reconstructs and measures DBH from above-canopy UAV RGB imagery by using a 3D Gaussian Splatting scene representation in a pixel-limited regime.
@@ 2 Materials
Data were collected in a managed Pinus taeda (loblolly pine) stand in southeastern Arkansas, USA. We established ten 0.2-acre circular plots (plots 1–10; radius 16.05 m; Fig. 3). At each plot center, we marked a circular plot boundary with a 16.05 m radius and measured all stems within the boundary. Each included tree was assigned a unique ID. For each tree, we recorded: (i) bearing and distance from the plot center; (ii) DBH measured with a diameter tape at 1.37 m above ground level; and (iii) height measured using a laser rangefinder.
@@
For the full set of plots, we collected overlapping RGB imagery using a DJI Matrice 4E in two campaigns. Imagery was captured with high overlap and mixed viewing geometry (including oblique views at 70 and 90 m altitude) to support SfM/MVS reconstruction and increase the chance of observing trunk surfaces through gaps in the canopy and understory. To improve absolute positioning consistency across flights, we corrected the onboard UAV GNSS trajectory via post-processed kinematic (PPK) processing using reference data from a nearby CORS station (ARMO), and used the corrected solution to update image geotags before reconstruction. An example UAV RGB flight trajectory is shown in Figure 4(a).
Figure 5: TreeDGS pipeline. RGB images are reconstructed with SfM/MVS and represented as 3D Gaussian splats $G$. We surface-sample dense points using opacity $\alpha$ and weights $w$, segment trunks $T$, and estimate DBH by opacity-weighted circle fitting on trunk slices.
@@ 3.1 Problem Statement
where $\mathbf{P}_{i}$ are the calibrated camera matrices, $\mathcal{X}_{\text{dense}}$ is a densified point set from OpenMVS, $\mathcal{G}$ is the optimized Gaussian field, and $\mathcal{T}_{t}$ denotes trunk points for tree instance $t$.
@@ 3.2 Structure-from-Motion and Multi-View Stereo
SfM takes the matched correspondences across overlapping image pairs and estimates: (i) camera poses $\{\mathbf{P}_{i}\}$ and (ii) a sparse 3D point set. We use standard robust estimation and bundle adjustment from GLOMAP [37] to refine the reconstruction. We denote the resulting calibrated cameras as:
@@
These cameras are used both for Gaussian training (Sec. 3.3) and for the multi-view opacity tests during surface sampling (Sec. 3.4).
@@
Figure 6: Comparison of back-projection points with multi-view depth consistency and the proposed sample points. (a) Depth-fusion baseline: rendered depth points are back-projected into 3D and kept only if they pass multi-view depth-consistency checks. This produces large sheet-like artifacts near the ground/understory and incomplete stem surfaces (red dashed boxes), despite appearing consistent across views. (b) Our opacity-guided surface sampling with depth-aware point-wise compositing gives a stem geometry suitable for DBH estimation.
@@ 3.3 Reconstruction with Gaussian Splats
Given calibrated cameras $\{\mathbf{P}_{i}\}$ and an MVS-densified initialization $\mathcal{X}_{\text{dense}}$ (Sec. 3.2), we optimize a 3D Gaussian Splatting (3DGS) scene representation. We adopt RaDe-GS [50], which builds on the real-time 3DGS renderer [22] while explicitly rasterizing depth in a way that is also useful for our subsequent surface-consistent point sampling.
@@
We integrate RLGS [28] as an online controller that adapts selected training hyper-parameters based on observed optimization dynamics. In our setting, this is used to stabilize opacity and geometry quality for the downstream DBH pipeline. After convergence, the optimized Gaussian field is denoted $\mathcal{G}$ and serves as the source representation for dense point sampling.
@@ 3.4 Surface Sampling with Opacity
Our downstream stages (3D trunk segmentation and DBH fitting) require a point cloud that is (i) dense enough to represent thin stems at stand-off distance and (ii) geometrically trustworthy under heavy occlusion from foliage and understory. This is not automatic from a 3D Gaussian Splatting (3DGS) reconstruction: the optimized field $\mathcal{G}$ is a collection of overlapping volumetric primitives whose parameters are trained through front-to-back alpha compositing rather than an explicit surface loss [22, 50]. As a result, two seemingly straightforward exports are unreliable in our UAV forest regime. First, exporting one point per Gaussian (the means) produces a fragmented and extremely sparse cloud (Sec. 4, Fig. 10), which is insufficient for robust 3D segmentation and measurement. Second, back-projecting rendered depth maps into 3D and fusing them via multi-view depth-consistency preserves view-consistent sheet-like artifacts from semi-transparent vegetation while eroding already-scarce trunk evidence, especially when stems are only a few pixels wide or intermittently visible (Fig. 6). Therefore, we sample densely in 3D, but we accept and score samples using tests that are consistent with the same compositing model that governs 3DGS rendering and optimization.
@@
The per-pixel accumulated alpha mask $A_{v}(\mathbf{u})$ tells us whether a ray intersects reconstructed mass somewhere along the ray, but it does not tell us whether a particular 3D sample $\mathbf{p}$ lies on the visible surface, behind it, or in front of it (all can project to the same pixel). For surface extraction, we need a depth-aware test that measures how much opacity accumulates in front of $\mathbf{p}$ along the viewing ray, using the same compositing model that governs training. RaDe-GS provides exactly this query by integrating point-wise alphas with a depth-plane formulation.
@@
The $\min(\cdot)$ depth clamping in Eq. (6) is critical for surface consistency: once $\mathbf{p}$ lies behind the locally rasterized surface depth $d_{k,v}(\cdot)$, the depth residual term saturates rather than increases with $r_{v}(\mathbf{p})$, preventing spurious “volumetric” accumulation behind the surface. RaDe-GS composites these point-wise alphas using the same front-to-back transmittance as in Eq. (5), yielding the cumulative opacity
@@
Then, we store two per-point reliability signals:
@@ 3.5 Semantic Trunk Extraction
The sampled point cloud from Section 3.4 contains trunks mixed with foliage, understory vegetation, and occasional floating samples caused by Gaussians with nonzero opacity in free space. To isolate stem geometry for DBH measurement, we used ForestFormer3D [48], a 3D semantic segmentation model, to predict a class label for each sampled point and to assign per-tree instance IDs when available. In our pipeline, the per-point opacity reliability $\bar{\alpha}$ is kept as an auxiliary output feature alongside $(x,y,z)$. We retain only points predicted as trunk to form a trunk-only cloud for each tree instance $t$, denoted $\mathcal{T}_{t}=\{(\mathbf{x}_{k},\bar{\alpha}_{k})\}$. This separation step reduces contamination from branches and understory, while the retained opacity values enable reliability-weighted fitting downstream. Then, the trunk-only instances are passed to the DBH measurement stage.
@@ 3.6 Opacity-Weighted DBH Measurement
Even after per-slice robust fitting, some slices can be corrupted by branch attachments, residual foliage, or incomplete sampling. Therefore, we fit a simple taper model on the lower stem using RANSAC with a negative-slope prior:
@@
Table 2: Per-plot DBH accuracy and robustness. RMSE/ MAE/ME are in cm and RRMSE is in %. SR reports successful estimates over total trees. Bold indicates the best within each plot. TreeDGS with opacity-weighted circle fitting attains the best RMSE in 7/10 plots and the best aggregated performance.